OpenStudy (anonymous):

4. A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

since this involves absolute value then we'll have: \[\left| x - 0.525 \right| = 0.01/-0.01\] note that when talk of absolute value, all the numbers are positive therefore we will have two answers. CASE 1: when it is equal to 0.01 removing the absolute value sign x - 0.525 = 0.01 x = 0.01 + 0.525 x = 0.535 CASE 2: when it is equal to -0.01 removing the absolute value sign x - 0.525 = -0.01 x = -0.01 + 0.525 x = 0.515